Solve by process of elimination:
n+d+q=75
.05n+.10d+0.75q=7.25
n=5d
*this question is dealing with coins*

Answer by checkley77(12844)

n+d+q=75
.05n+.10d+[0.75]q=7.25 should be .25q.
n=5d
5d+d+q=75
6d+q=75
----------
.05(5d)+.10d+.25q=7.25
.25d+.10d+.25q=7.25
.35d+.25q=7.25 multiply by -4 & add.
---------------------
-1.40d-q=-29
6d+q=75
------------
4.6d=46
d=46/4.6
d=10 dimes
n=5*10=50 nickrls.
10+50+q=75
60+q=75
q=75-60
q=15 quarters.
Proof:
.05*50+.10*10+.25*15=7.25
2.50+1.00+3.75=7.25
7.25=7.25